OpenStudy (anonymous):

What is the degree measure of an arc 4 ft. long in a circle of radius 10 ft.? 72° 60° 57.6°

OpenStudy (anonymous):

@aric200

OpenStudy (aric200):

I actually know I have seen this one in the past Find C C=2 * pi * r 2 * 3.14 * 10 = 62.8 360 * 4 = 1440 1440 * 3.14 = 4521.6 4521.6 / 62.8 = 72 @ZhangYan http://openstudy.com/study#/updates/4fc8f977e4b0c6963ad34e40

OpenStudy (anonymous):

wow thanks what kinda school work are you working on?

OpenStudy (aric200):

Basic stuff, Nonlinear Systems

OpenStudy (anonymous):

oh

OpenStudy (aric200):

I just have so much school It time consuming

OpenStudy (anonymous):

yeah. i know exactly what you mean. are you in Cali?

OpenStudy (aric200):

Right Now I am in Washington. I lived there a very short amount of town, but I'm always all over the place. I live In Arizona But moved from Florida to Arizona. I travel a lot, move a lot. I love California though

OpenStudy (anonymous):

oh dang. well i have all these assignments due by tomorrow and i live in georgia so its 3 here

OpenStudy (aric200):

My school is done Thursday So I know what you mean. Good Luck Ill try to help you as much as I can but if anything just keep at it ask if you need to hopefully people will help if I can not.

OpenStudy (anonymous):

thanks man